OpenStudy (anonymous):

PLEASE HELP ME:(on the unit cirlce what are the coordinates of the terminal point determined by theta =19pi/3

OpenStudy (anonymous):

A.(-1/2,-\[\sqrt{3}\div2\] ) B.(1/2,-\[\sqrt{3}\div2\] ) C.(-1/2,\[\sqrt{3}\div2\] ) D.(1/2,\[\sqrt{3}\div2\] )

OpenStudy (radar):